OpenStudy (anonymous):

A base following basket contains the following pieces of fruit: 3 apples, 2 oranges, 2 bananas, 2 pears, and 5 peaches. Jameson picks a fruit at random and does not replace it. Then Brittany picks a fruit a random. What is probability that Jameson gets a banana and Brittany gets a pear? a.4/27 b.1/49 c.2/91 d.27/91

OpenStudy (anonymous):

Chance banana first pick = 2 / 14 Chance pear second pick = 2 / 13(one is already picked) Because it states Jameson gets a banana AND Brittany gets a pear you need to multiply, if it was OR you need to add So 2/14 x 2 / 13 = 0.02178 = 2/91

OpenStudy (anonymous):

thank you @PatN